OpenStudy (anonymous):

Evaluate the expression -12 - 6a - 4b for each a and b variable pair. Match the variable pair with the resulting value of the expression. Match Term Definition ? a=1 & b=2 A) -14 ? a=-1 & b=-2 B) -26 ? a=-1 & b=2 C) -10 ? a=1 & b=-2 D) 2 Please help, I need to check these.

OpenStudy (anonymous):

Just replace, in expression " -12 -6a -b", a and b by those given in each line : e.g. : For a=1 & b = 2 ==> -12-6*1-4*2 = -12-6-8=-26 ==> MATCH definition B